OpenStudy (anonymous):

Let f(x) = 1/2(4)^x Find x when f(x)=32 It says the answer is 3???

OpenStudy (precal):

\[f(x)=\frac{1}{2}4^x\] Is this the problem?

OpenStudy (anonymous):

Yes, and then it asks to solve when f(x)=32

OpenStudy (precal):

\[32=\frac{1}{2}4^x\] multiply both sides by 2

OpenStudy (precal):

\[64=4^x\] yes 4 to the third power is 64

OpenStudy (precal):

Have you been taught about common logs or natural logs?

OpenStudy (anonymous):

Ah....I was dividing both side by 1/2 rather than multiplying :(

OpenStudy (anonymous):

Yes...when I logged it out I would comeup with two. Its because I was ending up with 16 rather then 64

OpenStudy (precal):

Yes, you can divide both sides by (1/2) but it would be easier to multiply by its reciprocal

OpenStudy (precal):